OpenStudy (anonymous):

Find dy/dx: y= (sin(x))^-5 - (cos(x))^3

OpenStudy (anonymous):

We have: \[y=\sin^5(x)-\cos^3(x)\] So then we need to apply the chain rule which is: \(\frac{d}{dx}(f\circ g)(x)=f'[g(x)]g'(x)\) We obtain: \[\eqalign{ \frac{dy}{dx}&=5\sin^4(x)\cos(x)-3\cos^2(x)(-\sin(x)) \\ &=5\sin^4(x)\cos(x)+3\cos^2(x)\sin(x) \\ }\]

OpenStudy (anonymous):

Oh I get it now thanks :)

OpenStudy (anonymous):

Anytime :)